OpenStudy (anonymous):

Solve: 2x + 36 + 40 = 180 52 128 88 92

OpenStudy (anonymous):

36 + 40 =

OpenStudy (anonymous):

x=52

OpenStudy (anonymous):

the answer is 52

OpenStudy (andriod09):

you have the equation: \[2x+36+40=180\]\[2x+76=180\]\[-76=-76\]\[2x=104\]\[x=52\]